OpenStudy (anonymous):

A new software company wants to start selling DVDs with their product. The manager notices that when the price for a DVD is 16 dollars, the company sells 150 units per week. When the price is 27 dollars, the number of DVDs sold decreases to 82 units per week. Answer the following questions: (A) Assume that the demand curve is linear. Find the demand, q, as a function of price, p. (B) Write the revenue function, as a function of price. (C) Find the maximum revenue.

OpenStudy (anonymous):

@amistre64 Hi! Can you help me again?

OpenStudy (amistre64):

you have 2 points of reference, right?

OpenStudy (anonymous):

(16,150) (27,82)

OpenStudy (anonymous):

When I calculated the slope, I got -68/11

OpenStudy (amistre64):

thats what the slope if between the points yes

OpenStudy (amistre64):

f(p) = slope(p-po) + f(po)

OpenStudy (anonymous):

I thought we were supposed to use q= mp + b ?

OpenStudy (amistre64):

there is more than one form of a line, but if you must m(p-po) + f(po) can be distributed and collected as mp+b

OpenStudy (anonymous):

Well I ended up getting q=(-68/11)p + 249 but it said it is incorrect :/

OpenStudy (anonymous):

And I don't know what I did wrong

OpenStudy (amistre64):

-68/11(x-27)+82 2738/11-(68 x)/11 id say you mathed an error into it

OpenStudy (amistre64):

if we know y=mx+b, then b=y-mx right? 82 + 68/11 (27) = b

OpenStudy (anonymous):

248.9

OpenStudy (anonymous):

But that's the same thing as 249 haha

OpenStudy (amistre64):

.909090... but yeah :)

OpenStudy (amistre64):

now how do we determine revenue?

OpenStudy (anonymous):

wait so why is q=(-68/11)p + 249 wrong? :/

OpenStudy (amistre64):

becasue 249 is not right

OpenStudy (amistre64):

keep it as a faction

OpenStudy (amistre64):

or fraction even

OpenStudy (amistre64):

q = (-68p+2738)/11

OpenStudy (amistre64):

or q = -(68p-2738)/11 either way to me

OpenStudy (amistre64):

revenue is just price times quantity sold ... or pq

OpenStudy (anonymous):

Ok got it! I just need help with part D :)

OpenStudy (amistre64):

youll need to post part D

OpenStudy (anonymous):

Oops sorry C

OpenStudy (amistre64):

what is your quadratic from B?

OpenStudy (anonymous):

q= -68/11p^2+248.9p

OpenStudy (amistre64):

if you havent learned how to determine the min or max of a U shaped curve by now .. theres not much i can do to save you lol

OpenStudy (anonymous):

Do you have to graph it?

OpenStudy (amistre64):

you can, if you must

OpenStudy (amistre64):

the min or max of a quadratic occurs at the middle of the roots ....

OpenStudy (amistre64):

pq has roots when p=0, or q=0

OpenStudy (amistre64):

what is the q intercept? then half it

OpenStudy (amistre64):

or just do -b/(2a)

OpenStudy (amistre64):

well, that is not the max, but that is where the max occurs

OpenStudy (anonymous):

I got -20.13 but that's wrong

OpenStudy (amistre64):

20.13 is the price that gets us the most revenue ... let p=20.13 ... what is the revenue from it?

OpenStudy (amistre64):

strategy: define the line q=mp+b from the points of reference define the revenue as pq determine the value of p, that gets us the most revenue use that value to determine the max revenue